### v4

@@ 4.5 Angular relative positional encoding
Since follows a radial partitioning, we employ an angular relative positional encoding to capture the relative position between tokens with respect to incident $\theta$ and azimuth $\varphi$ angles. We divide the position bias $B$ in eq. 6 into two parts: $B_{\theta}$ and $B_{\varphi}$, the incident-angular and azimuthal relative position bias respectively. Given the $i$-th token angular coordinates
@@
where $\theta_{\mathrm{max}}$ is the half field of view. The relative angular positions $(\Delta\theta,\Delta\varphi)$ between tokens $i,j$ are
@@
Here, $a_{*}$ and $b_{*}$ are trainable coefficients. Since the relative positions in a window along angular and azimuth axes ranges from $[-M_{\theta}+1,M_{\theta}-1]$ and $[-M_{\varphi}+1,M_{\varphi}-1]$ respectively, where $M_{\theta}$ and $M_{\varphi}$ are number of patches in a window (see sec. 4.4, here $M_{\theta}=M_{r}$). We parameterize two bias matrices $\hat{B}_{\varphi}\in\mathbb{R}^{(2M_{\varphi}-1)\times 2}$ and $\hat{B}_{\theta}\in\mathrm{R}^{(2M_{\theta}-1)\times 2}$. Hence $a_{*}$ and $b_{*}$ for a particular token are sampled from $\hat{B}_{\theta}$ and $\hat{B}_{\varphi}$ based on the relative positions of the said token to generate angular and azimuth position values in $B_{\theta}$ and $B_{\varphi}$.
@@ 4.6 Polar patch merging
Similar to window-based self attention (sec. 4.4), the polar nature of our architecture enables many possibilities when merging patches. For example, we could merge $2\times 2$ neighboring patches (-RA) or $1\times 4$ merge along azimuth (-A) as shown in tab. 1. We found the azimuth merging strategy to outperform the others in our experiments (see sec. 5.6).
4.7 Polar patch expanding
As in [27], we use an MLP layer for the expanding layer. We use this layer along the azimuth dimension to upsample by a 4×4$\times$4 × factor. For example, consider the first patch expanding layer in fig. 2. The input feature map $(N_{r}\times\frac{N_{\varphi}}{64}\times 8c)$ is first given to an MLP layer to expand the feature dimension by 4×4$\times$4 × to get $(N_{r}\times\frac{N_{\varphi}}{64}\times 32c$). The feature map is then rearranged to reduce the feature dimension and increase the resolution of feature map along azimuth dimension to obtain $(N_{r}\times\frac{N_{\varphi}}{16}\times 4c$).
4.8 $k$-NN layer
Lastly, we employ a $k$-NN layer to map the polar feature map back to cartesian coordinates. Each pixel coordinate in the image is associated with its $k$ closest samples (we use $k=4$), and their respective feature vectors are averaged. Since sample point locations are known, the $k$-NN layer is fixed and not trainable. The $k$-NN output (of dimensions $H\times W\times c$ is fed into the last linear projection layer to get the desired output for the required task.
4.9 Data availability
All data supporting the findings of this study are available within the paper and explained in sec. 5.1 and sec. 6.1.
Table 1: Model architecture specification and variants. In both variants, we extend the architecture of DarSwin [26] to DarSwin-Unet, inspired by Swin-Unet [27], which has three stages of patch merging, two stages of bottleneck, and three stages of patch expanding layer. “Resolution” : resolution of input feature map at every stage. “Window Size” ($M_{r},M_{\varphi}$): Number of patches along radial $M_{r}$ or azimuth $M_{\varphi}$ axis in a window. “-RA”: encoder architecture variant merging $2\times 2$ neighborhood patches along the radius and azimuth. “-A”: encoder architecture variant merging $1\times 4$ neighborhood patches along the radius and azimuth. “$k$-NN”: layer to map the output feature map from polar back to cartesian coordinates.
-RA -A
Resolution
@@ 4.9 Data availability
Stage-4
(2, 8)
(2, 8)
Stage-4
(16, 1)
(16, 1)
Stage-3
(4, 16)
(4, 4)
Stage-3
(16, 4)
(4, 4)
Stage-2
(8, 32)
(4, 4)
Stage-2
(16, 16)
(1, 16)
Stage-1
(16, 64)
(4, 4)
Stage-1
(16, 64)
(1, 16)
$k$-NN
(64, 64)
N/A
$k$-NN
(64, 64)
N/A
@@ 5 Classification experiments
We begin by evaluating the efficacy of our proposed encoder only (i.e., top row of fig. 2). For this, we perform a series of experiments on image classification. Since there exists no classification dataset for wide-angle images, we instead create synthetically distorted images using 200 randomly chosen classes from the ImageNet1k dataset [48].
Figure 6: Mean top-1 classification accuracy as a function of input lens distortion for our method (-A) and previous state of the art: DAT [25], Swin [5], and Swin(undis) (see text). All methods are trained on a restricted set of lens distortion curves (indicated by the pink shaded regions), from top-left: very low, low, medium and high distortion. We observe zero-shot adaptation to lens distortion of each approach by testing across all $\xi\in[0,1]$.
@@ 5.1 Dataset
Test sets of 30,000 images over 200 classes are generated using the same procedure. Here, $\xi$ is determined once for each image and kept fixed. We test for all $\xi$ values between $[0,1]$ to evaluate generalization to different lens distortions.
@@ 5.2 Baselines and training details
We perform classification experiments using architecture, which can be obtained by appending a 200-unit linear layer to the output of the stage 4 block of the encoder in fig. 2. We compare our approach with the following baselines: Swin-S [5], Deformable Attention Transformer (DAT-S) [25], and Swin-S on input images that have been undistorted back to a perspective projection (dubbed “Swin(undis)”). As with our , this last baseline has knowledge of the distortion parameters, whereas the first two do not. Note that we do not include comparisons to methods that estimate distortion [14, 15, 41, 11]. Indeed, the spherical projection model (sec. 3) is bijective: the undistortion function has a closed form and is exact. Therefore, the Swin (undis) method serves as an upper bound to all self-calibration methods because it is, in essence, being given the “ground truth” undistortion.
@@
All three baselines employ 32 divisions along the image width and height. For , we use $N_{r}=16$ and $N_{\varphi}=64$ divisions along the radius and azimuth respectively, which yields the same total number of 1024 patches for all methods. All three baselines are trained with a window size (4, 4) on our synthetically distorted training sets.
@@
All methods use the AdamW optimizer on a batch size of 128 using a cosine decay learning rate scheduler and 20 epochs of linear warm-up. We use an initial learning rate of 0.001 and a weight decay of 0.05. We include all of the augmentation and regularization strategies of [5], except for random crop and geometric transformation (like shearing and translation). Our model requires 0.03M additional parameters over the Swin baseline, which contains 48M, representing a 0.061% increase.
Figure 8: Sensitivity analysis to errors in lens distortion parameter $\xi$. Each plot shows the color-coded mean top-1 classification accuracy (brighter is better). For each plot, the $x$ and $y$ axes represent the error $\Delta\xi$ and the ground truth in distortion parameter $\xi$ respectively given to (top row) and Swin(undis) (bottom row) during evaluation. Columns correspond to different distortion levels (from Very Low to High) used during training.
Figure 9: Results for the self-calibration experiments, where the lens distortion parameter $\xi$ is provided, at test time, by a separate calibration network. We compare three training scenarios for : trained on ground truth $\xi$ (“pre-trained”), trained on ground truth $\xi$ then fine-tuned on predictions from the calibration network (“fine-tuned”), and trained from scratch on predictions from the calibration network (“from scratch”).
5.3 Zero-shot generalization
@@
We are interested in evaluating whether our distortion-aware method can better generalize to other unseen lenses at test time. Hence, we train each approach on a specific distortion level (c.f. sec. 5.1) and evaluate them on all distortion values $\xi\in[0,1]$. For this, the entire test set is distorted using a single $\xi$ value, and we repeat this process for every $\xi$ in uniform discrete interval $[0,1]$ with step size 10, to simulate different lens distortion.
@@
Results are reported in fig. 6, where the distribution of training distortions is drawn in pink. We observe that performs on par with the baselines when test distortions overlap with training, but shows much greater zero-shot generalization capabilities outside of the training lens distortion domain. Furthermore, tab. 2 shows the top-1 accuracy for the test set distorted with $\xi=0.4$, which is in none of the training intervals. Again, we note that yields better generalization accuracy (without fine-tuning) than all baselines, even Swin(undis) which also has access to the ground truth distortion function (see sec. 5.2).
@@
Swin (undis) 47.48 91.52 91.07 87.34
DAT 57.5 90.4 88.5 75.7
Table 2: Comparisons of mean top-1 accuracy of the models trained on four levels of distortion (pink regions in fig. 6) and tested on distortion level $\xi=0.4$. Each row is color-coded as best and second best.
5.4 Sensitivity analysis
Recall that all experiments so far assume knowledge of the ground truth lens distortion parameter, $\xi$, of each image at both training and test time. We are interested in evaluating the sensitivity of approaches which assume knowledge of the lens distortion curve (namely, and Swin(undis)) to errors in $\xi$ at test time. To this end, we inject an error $\Delta\xi\in[-0.4,0.4]$ during the evaluation of trained models and show the results in fig. 8. As can be observed, is overall much more robust to errors in the distortion parameter than Swin(undis). This is especially visible when models are trained on the “very low” group. In this scenario, the accuracy of Swin(undis) severely drops when a small error in $\xi$ is provided to the network. In contrast, the performance of remains much more stable across even large values of $\Delta\xi$.
5.5 Self-calibration
Given the robustness of to errors in lens distortion curve, we also experiment with combining it with a separate “calibration” network whose task is to estimate the lens distortion parameter $\xi$ from the input image itself.
Training a calibration network
Here, we follow the same approach as [72] and employ a ConvNext [73] encoder and replace the last layer with a linear layer which predicts the distortion $\xi$ directly. The output layer uses the softmax activation function on 256 outputs, which correspond to a uniform binning of the $[0,1]$ interval. We employ the cross entropy loss on all outputs and train the model using the Adam optimizer. We train this network on the ImageNet1k train set by distorting the images on the fly through uniform sampling $\xi\in[0,1]$.
Self-calibration experiments
We perform three sets of experiments and report results in fig. 9. In all cases, is provided with the lens distortion parameter $\xi$ predicted by the calibration network at test time. We experiment with three training scenarios where is trained: 1) on ground truth $\xi$ exclusively; 2) on ground truth $\xi$ then fine-tuned on predictions from the calibration network; and 3) from scratch on predictions from the calibration network.
We observe in fig. 9 that while all scenarios perform similarly within the training region, both the “fine-tuned” and “from scratch” markedly better in generalization. Indeed, adapting to the predictions of the calibration network helps, as expected. There does not seem to be a noticeable difference between “fine-tuned” and “from scratch”, however. This last result shows that can be trained on uncalibrated wide-angle images, provided a robust calibration network is available (such as [72] for example).
@@ 5.6 Ablations
We ablate some important design elements (tab. 3) and training strategies (tab. 4). All ablations are performed using -A trained on the “low” distortion ($\xi\in[0.2,0.35]$) and tested on all the distortion levels.
@@
We compare our angular relative positional encoding (sec. 4.5) with two versions of the Fourier-based polar positional encoding from [74] in the first part of tab. 3. First, polar PE encodes patches using their relative distortion-aware radial lengths $r$ and azimuth values $\varphi$. Second, angular PE encodes patches using the relative value of incident angle $\theta$ and azimuth $\varphi$. While angular PE yields better performance, we observe that our angular relative PE improves the performance even further.
Number of sampling points along radius $N_{sr}$
From $N_{s}$ sampling points in and $N_{sr}$ defining number of sampling points along radius. We observe in tab. 3 that, with $N_{sr}=10$ sampling points along the radius outperforms the models with $N_{sr}=5$ and $N_{sr}=2$ sampling points along the radius by just over (1-6)%.
@@
Table 3: Ablation study on different design elements of , including different positional encodings, number of sampling points along the radius $S_{r}$, and window formation and merging strategy: -A (azimuth) or -RA (radius+azimuth).
Jittering
When patches are sampled for the linear embedding layer (c.f. sec. 4.3), an augmentation strategy is used to jitter the sample points in the patch. According to tab. 4, this jittering augmentation improves the performance by (1–2)%.
Table 4: Ablation study on different types of sampling techniques and augmentation strategies on -A. For jittering we compare “no jitter” (all samples are given to the MLP for linear embedding as is) with “with jitter” (jittering is applied). For sampling, we compare “Unif. sampl.”: points are sampled uniformly inside a patch; to “Distortion aware (DA) sampl.”: lens information is taken into account to sample points inside a patch. ”Distortion aware (DA) part.” uses lens information to partition the patch.
@@
With jitter 83.4% 92.8% 91.5% 88.3%
No jitter 81.5% 92.6% 91.2% 87.4 %
DA sampl. 83.4% 92.8% 91.5% 88.3%
Unif. sampl. 82.4% 92.5% 91.2% 86.2%
DA part. 83.4% 92.8% 91.5% 88.3%
w/o DA part. 52.2% 91.3% 86.2% 75.5%
@@
We observe in tab. 4 that without lens information, cannot generalize to unseen distortion.
@@ 5.7 Generalization over projection models
In tab. 5, we aim to check for robustness to domain shift at inference due to the use of a different distortion model (c.f. sec. 3). In particular, we use the test set of 30,000 images and distort them using the distortion parameters of a 4-degree polynomial projection model. Each image in the test set is assigned four different distortion parameters randomly sampled from a uniform distribution. The undistortion is not exact since the polynomial projection function is not invertible. Hence Swin(undis) performs only better than on low distortions, but the performance degrades on medium to high distortion levels. Our model performs better generalization on all distortion levels. DAT fails to generalize on low distortion levels, and Swin (undis) fails to generalize on high distortions.
@@
Very Low Low Medium High
Swin(undis) 85.9% 87.1% 71.3% 60.3%
DAT 65.4% 85.9% 78.9% 63.8%
@@
82.7% 84.8% 78.4% 65.5%
Figure 10: From an input RGB panorama (top) with corresponding depth map (not shown) from the Matterport3D dataset [75], we generate a wide angle image and its depth map (bottom). Here, the field of view is ${}^{\circ}$ and $\xi=0.95$.
(a) Wide angle image
(b) Partial cubemap
Figure 11: Cubemap generation: we warp the (a) $175^{\circ}$ wide angle image to the corresponding (b) partial cubemap representation.
6 Depth estimation experiments
We now experiment with our full distortion-aware encoder-decoder architecture (DarSwin-Unet) from fig. 2. Monocular depth estimation is considered as target task for this purpose, since it is a pixel-level regression problem that is inherently ill-posed and ambiguous because many 3D scenes can be projected to the same 2D image. The problem also becomes more complex with the significant distortion in fisheye images. Experiments provided in this section confirm the previous findings about ’s promising generalization abilities and show that they apply for pixel-level tasks as well. In this section, we will refer to our method as DarSwin-Unet to emphasize its encoder-decoder structure.
Figure 12: Depth estimation accuracy $\delta_{1}$ (higher is better) as a function of test distortion for: DarSwin-Unet, Swin-Unet [27], Swin(undis)-Unet, Swin-UPerNet [5, 76], Swin(undis)-UPerNet and DAT-UPerNet [25, 76]. All methods are trained on a restricted set of lens distortion curves (indicated by the pink shaded regions): (a) Very low, (b) low, (c) medium and (d) high distortion. We study the generalization abilities of each model by testing across all $\xi\in[0,1]$.
6.1 Datasets
Since existing wide-angle depth estimation datasets such as Woodscapes [4] do not provide a wide range of distortion profiles to understand the generalization capacity of our network, we instead simulate a wide range of lenses using the same spherical model (c.f. sec. 3) on the indoor Matterport3D dataset [75] as illustrated in fig. 10. Note that this warping can easily be done since the spherical projection model (c.f. sec. 3) allows a closed-form solution to its inverse projection.
Training set
Similar to the classification experiments (sec. 5.1), we generate four different training sets with the same levels of distortion. Training images are synthetically generated from panoramas on the fly during training with distortion $\xi$ sampled from their respective intervals and the yaw viewing angle being uniformly sampled in the $[0,360^{\circ}]$ interval. Each of the four training set (one for each distortion group) contains 9180 panoramas of original resolution of $512\times 1024$, where the generated wide-angle image is subsequently downsampled to $64\times 64$ after warping.
Test set
To evaluate zero-shot generalization to lens distortions, we generate wide-angle images from the test set containing 1620 panoramas, for every $\xi$ in uniform discrete distribution $[0,1]$ with step size 10.
6.2 Baselines and training details
We compare DarSwin-Unet with the following baselines: Swin-Unet [27], Swin-UPerNet (a Swin [5] encoder with UPerNet [76] decoder) and DAT-UPerNet (a DAT [25] encoder with UPerNet decoder). As with classification (sec. 5), we also wish to compare against models trained on undistorted images. Here, we include two such additional baselines: Swin-Unet(undis) and Swin-UPerNet(undis).
Undistortion-piecewise linear correction
Undistorting a $175^{\circ}$ field-of-view wide angle image to a single perspective image will result in extremely severe stretching. Instead, we follow the piecewise linear correction strategy in [4] and undistort the image to a partial cubemap, which is composed of 6 perspective faces of $90^{\circ}$ field of view each, unrolled into an image. As shown in fig. 11, this preserves the entire field of view while minimizing stretching. These images are used to train the “(undis)” baselines mentioned above.
Training details
All baselines employ 32 divisions along the image width and height. For DarSwin-Unet, we use $N_{r}=16$ and $N_{\varphi}=64$ divisions along the radius and azimuth respectively, which yields the same number of 4096 patches for image size $64\times 64$. All three baselines are trained with a window size $(4,4)$ and DarSwin-Unet is trained with window size $(1,16)$.
All encoders (Swin and DAT) are first pre-trained for classification on the distorted ImageNet1K dataset from sec. 5.1. Pre-trained encoders are then fine-tuned, along with their respective decoders, on the depth estimation task.
All methods are trained with the SGD optimizer with momentum $0.9$ and weight decay 1e-4 on a batch size of 8. We employ a polynomial learning rate policy with a base learning rate of $0.01$ and $\text{power}=0.9$. We use random flips and rotations as data augmentation.
Training loss
Since the global scale of a scene is a fundamental ambiguity in depth estimation [77], we train the network using the scale-invariant loss in log space [78]:
$L=\sqrt{\frac{1}{n}\sum_{i}{d_{i}}^{2}-\frac{\lambda}{n^{2}}(\sum_{i}{d_{i}})^{2}}\,,$ (12)
with $d_{i}$ is the difference between the predicted and ground truth (log-)depth. We use $\lambda=0.85$.
Evaluation metrics
We evaluate performance on typical depth estimation metrics [78]: absolute relative error, RMSE, log-RMSE, squared relative error and accuracy under threshold ($\delta_{i},\,i\in\{1,2,3\}$). The paper reports results on
$\delta_{1}=\frac{1}{|D|}|\{d\in D|\max(\frac{d^{*}}{d},\frac{d}{d^{*}})\leq 1.25\}\,,$ (13)
where $D$, $d^{*}$ and $d$ are the set of valid, ground truth and predicted depths respectively. Please consult the supplementary material for other metrics.
6.3 Zero-shot generalization
As with the classification experiments (sec. 5.3), we train each approach on a specific distortion level (c.f. sec. 5.1) and evaluate them on all distortion values $\xi\in[0,1]$. Fig. 12 illustrates the results using the accuracy $\delta_{1}$ metric. Overall, we observe that DarSwin-Unet maintains the most stable performance as we move away from the training distortion for all levels of training distortions. We note that Swin(undis)-Unet has competitive performance when the test distortion is greater than the one used for training. However, its performance severely drops in the opposite scenario (visible in the Low, Medium and High distortion figures). We also observe that the performance of DarSwin-Unet when the test distortion is equal to training (indicated by the shaded regions) is somewhat lower than other baselines (all cases except the Very Low training distortion): DAT-UperNet is the best-performing method in this case. However, its performance in generalization outside the training distortion severely drops, indicating that it cannot adapt to other lenses at test time.

### v5

@@ 4.5 Angular relative positional encoding
Since DarSwin follows a radial partitioning, we employ an angular relative positional encoding to capture the relative position between tokens with respect to incident $\theta$ and azimuth $\varphi$ angles. We divide the position bias $B$ in eq. 6 into two parts: $B_{\theta}$ and $B_{\varphi}$, the incident-angular and azimuthal relative position bias respectively. Given the $i$-th token angular coordinates
@@
where $\theta_{\mathrm{max}}$ is the half field of view. The relative angular positions $(\Delta\theta,\Delta\varphi)$ between tokens $i,j$ are given by
@@
Here, $a_{*}$ and $b_{*}$ are trainable coefficients. Since the relative positions in a window along angular and azimuth axes ranges from $[-M_{\theta}+1,M_{\theta}-1]$ and $[-M_{\varphi}+1,M_{\varphi}-1]$ respectively, where $M_{\theta}$ and $M_{\varphi}$ are number of patches in a window (see sec. 4.4, here $M_{\theta}=M_{r}$). We parameterize two bias matrices $\hat{B}_{\varphi}\in\mathbb{R}^{(2M_{\varphi}-1)\times 2}$ and $\hat{B}_{\theta}\in\mathrm{R}^{(2M_{\theta}-1)\times 2}$. Hence $a_{*}$ and $b_{*}$ are taken from $\hat{B}_{\theta}$ and $\hat{B}_{\varphi}$.
@@ 4.6 Polar patch merging
Similar to window-based self attention (sec. 4.4), the polar nature of our architecture enables many possibilities when merging patches. For example, we could merge $2\times 2$ neighboring patches (DarSwin-RA) or $1\times 4$ merge along azimuth (DarSwin-A) as shown in tab. 1. We found the azimuth merging strategy to outperform the others in our experiments (see sec. 5.4).
Table 1: Model architecture specification and variants. In both variants, we adapt the small architecture of Swin Transformer [28], which has four stages of patch merging. Input Resolution : Resolution of feature map at every stage. Window Size : Window size ($M_{r},M_{\varphi}$), $M_{r}$ : Number of patches along radial axis in a window, $M_{\varphi}$ : Number of patches along azimuth axis in a window. DarSwin-RA: Architecture variant with four patches along both azimuth and radial axis and merging $2\times 2$ neighborhood patches along the radius and azimuth. DarSwin-A: architecture variant with 16 patches along the azimuth axis and one patch along the radial axis in a window and merging $1\times 4$ neighborhood patches along the radius and azimuth.
DarSwin-RA DarSwin-A
Input Resolution
@@ 5 Classification experiments
To evaluate the efficacy of our proposed DarSwin encoder, we perform a series of experiments on image classification. Since there exists no classification dataset for wide angle images, we instead create synthetically distorted images using 200 randomly chosen classes from the ImageNet1k dataset [23].
@@ 5.1 Dataset
Test sets of 30,000 images over 200 classes are generated using the same procedure. Here, $\xi$ is determined once for each image and kept fixed. And we test for all $\xi$ values between $[0,1]$ to evaluate generalization to different lens distortions.
@@ 5.2 Baselines and training details
We compare our approach with the following baselines: Swin-S [28], Deformable Attention Transformer (DAT-S) [44], and Swin-S on input undistorted images dubbed “Swin(undis)”. As with our DarSwin, this last baseline has knowledge of the distortion parameters whereas the first two do not. Note that we do not include comparisons to methods which estimate distortion [47, 45, 21, 52]. Indeed, the spherical projection model (sec. 3) is bijective: the undistortion function has a closed form and is exact. Therefore, the “Swin (undis)” method serves as an upper bound to all self-calibration methods because it is, in essence, being given the “ground truth” undistortion.
@@
All three baselines employ 32 divisions along the image width and height. For DarSwin, we use $N_{r}=16$ and $N_{\varphi}=64$ divisions along the radius and azimuth respectively, which yields the same total number of 1024 patches for all methods. All three baselines are trained with a window size (4, 4) on our synthetically distorted training sets.
@@
All methods use the AdamW optimizer on a batch size of 128 using a cosine decay learning rate scheduler and 20 epochs of linear warm-up. We use an initial learning rate of 0.001 and a weight decay of 0.05. We include all of the augmentation and regularization strategies of [28], except for random crop and geometric transformation (like shearing and translation). Our model requires 0.03M additional parameters over the Swin baseline which contains 48M, representing a 0.061% increase.
Refer to caption      Refer to caption
(a) Very low (b) Low
Refer to caption      Refer to caption
(c) Medium (d) High
Figure 7: Top-1 classification accuracy (mean) as a function of test distortion for our method (DarSwin-A) and previous state of the art: DAT [44], Swin [28], and Swin (undis) (see text). All methods are trained on a restricted set of lens distortion curves (indicated by the pink shaded regions): (a) Very low, (b) low, (c) medium and (d) high distortion. We observe zero-shot adaptation to lens distortion of each approach by testing across all $\xi\in[0,1]$.
5.3 Zero-shot lens distortion generalization
@@
We are interested in evaluating whether our distortion-aware method can better generalize to other unseen lenses at test time. Hence, we train each approach on a level of distortion (c.f. sec. 5.1) and evaluate them on all distortion values $\xi\in[0,1]$. For this, entire test set is distorted using a single $\xi$ value, and we repeat this process for every $\xi\in[0,1]$ to simulate different lens distortion.
@@
Results are reported in fig. 7, where the distribution of training distortions is drawn in pink. We observe that DarSwin performs on par with the baselines when test distortions overlap with training, but shows much greater generalization capabilities outside of the training domain. Furthermore, tab. 2 shows the top-1 accuracy for the test set distorted with $\xi=0.4$, which is in none of the training intervals. Again, we note that DarSwin yields better generalization accuracy (without fine-tuning) than all baselines, even Swin (undis) which also has access to the ground truth distortion function (see sec. 5.2).
Table 2: Comparisons of top-1 accuracy of the models trained on four levels of distortion (pink regions in fig. 7) and tested on distortion level $\xi=0.4$. Each row is color-coded as best and second best.
@@
[…]
@@ 5.4 Ablations
We ablate some important design elements (tab. 3) and training strategies (tab. 4). All ablations are performed using DarSwin-A trained on the “low” distortion ($\xi\in[0.2,0.35]$) and tested on all the distortion levels.
@@
We compare our angular relative position encoding (sec. 4.5) with two versions of the Fourier-based polar positional encoding from [32] in the first part of tab. 3. First, polar PE encodes patches using their relative distortion-aware radial lengths $r$ and azimuth values $\varphi$. Second, angular PE encodes patches using the relative value of incident angle $\theta$ and azimuth $\varphi$. While angular PE yields better performance, we observe that our angular relative PE improves the performance even further.
Number of sampling points along radius $S_{r}$
We observe in tab. 3 that DarSwin with $S_{r}=10$ sampling points along the radius outperforms the models with $S_{r}=5$ and $S_{r}=2$ sampling points along the radius by just over (1-6)%.
@@
Table 3: Ablation study on different design elements of DarSwin, including different positional encoding, number of sampling points along the radius $n_{r}$, and window formation and merging strategy: DarSwin-A (azimuth) or DarSwin-RA (radius+azimuth ).
@@
Table 4: Ablation study on different types of sampling techniques and augmentation strategies on DarSwin-A. For jittering we compare “no jittering” (all samples are given to the MLP for linear embedding as is) with “with jittering”( jittering is applied on the sample points in a patch). For sampling, we compare “Uniform sampling”: points are sampled uniformly inside a patch; to “Distortion aware (DA) sampling”: lens information is taken into account to sample points inside a patch. "Distortion aware (DA) partition" DarSwin uses lens information to partition the patch.
@@
With jittering 83.43% 92.8% 91.5% 88.27%
No jittering 81.5% 92.6% 91.16% 87.4 %
DA sampling 83.43% 92.8% 91.5% 88.27%
Uniform sampling 82.4% 92.5% 91.2% 86.2%
DA partition 83.43% 92.8% 91.5% 88.27%
w/o DA partition 52.2% 91.3% 86.2% 75.5%
Jittering.
As discussed in sec. 4.3, when patches are sampled for the linear embedding layer, an augmentation strategy is used to jitter the sample points in the patch. According to tab. 4, this jittering augmentation improves the performance by (1–2)%.
@@
We observe in tab. 4 that without lens information, DarSwin cannot generalize to unseen distortion.
@@ 5.5 Generalization over projection models
In tab. 5, we aim to check for robustness for domain shift at inference due to the use of a different distortion model (c.f. sec. 3). In particular, we use the test set of 30,000 images and distort them using the distortion parameters of a 4-degree polynomial projection model. Each image in the test set is assigned four different distortion parameters randomly sampled from a uniform distribution. The undistortion is not exact since the polynomial projection function is not invertible. Hence Swin (undis) performs only better than DarSwin on low distortions, but the performance degrades on medium to high distortion levels. Our model performs better generalization on all distortion levels. DAT fails to generalize on low distortion levels, and Swin (undis) fails to generalize on high distortions.
@@
Training Levels Very Low Low Medium High
[…]
@@
[…]